Circular No. 7428

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

POSSIBLE RADIO SUPERNOVA IN NGC 6240 J. F. Gallimore, National Radio Astronomy Observatory, reports the

discovery of a possible radio supernova in the galaxy merger NGC 6240 in 2.3-GHz (13-cm) Very Long Baseline Array observations on 1999 Aug. 15 and 2000 Apr. 24. The two parent galaxy nuclei were detected at α = 16^{h}52^{m}58^{s}.9242, δ = +2^{o}24′04′′.771 and (end figures) 58^{s}.8902, 03′′.334 (equinox 2000.0; absolute astrometry accurate to 0′′.2; relative astrometry accurate to 0′′.001). The radio flux densities of the parent nuclei did not change between the two dates. The possible radio supernova, which dimmed from 1.3 to 0.5 mJy between the two dates, was detected at position end figures 58^{s}.8998, 03′′.591, which is 0′′.144 east and 0′′.257 north of the southern nucleus. The radio luminosity of this possible supernova is comparable to the very luminous radio supernovae detected in Arp 220 by Smith et al. (1998, Ap.J. 493, L17). Gallimore adds that only two nuclear sources were present in earlier MERLIN imagery, obtained by himself at 5 GHz on 1998 Jan. 12 (limit < 1 mJy, 3σ; 10-hr observation) and by R. Beswick, NRAL, Jodrell Bank, at 1.4 GHz on 1996 Jan. 31 (confusion-limited; flux density < 3 mJy). SN 2000bg (cf. IAUC 7392) is also in NGC 6240.

SUPERNOVA 2000cj IN NGC 6753 N. Suntzeff, Cerro Tololo Interamerican Observatory (CTIO), reports

that his reduction of a spectrum (range 450–670 nm, resolution 1900), obtained by A. Kundu (University of Virginia) on May 20.3 UT with the Las Campanas Swope 1-m telescope, shows that SN 2000cj (cf. IAUC 7425) is a type-Ia supernova very close to maximum light.

AX J0051.6–7311 K. Torii, National Space Development Agency of Japan; and J. Yoko-

gawa, K. Imanishi, and K. Koyama, Kyoto University, report that an ASCA observation made during Apr. 11–17 detected coherent pulsations of period 172.40 ± 0.03 s from a new source in the Small Magellanic Cloud, designated AX J0051.6–7311 and located at α = 0^{h}51^{m}37^{s}.8, δ = −73^{o}11′01′′ (equinox 2000.0; error radius 1′.5). The spectrum is characterized by a flat power-law function with photon index 0.9 and x-ray flux 1.4 × 10^{−12} erg cm^{−2} s^{−1} (0.7–10 keV). Within the error region of AX J0051.6–7311 is a ROSAT source known as RX J0051.9–7311 (Cowley et al. 1997, PASP 109, 21), having a Be star an an optical counterpart.

2000 May 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green